Calculate projected Net Liquidity in dollars for the Star Bank, N.A. from the following data: (*Remember,...

  1. Calculate projected Net Liquidity in dollars for the Star Bank, N.A. from the following data: (*Remember, this is all about cash availability)

  1. -deposit withdrawals will total $125 million;
  2. -customer loan repayments will amount to $88 million;
  3. -operating expenses in cash payment should be $51 million;
  4. -approved new loan requests total $205 million;
  5. -loan loss provision should increase by $10 million (*);
  6. -sales of securities should net $20 million;
  7. -new deposits should total $320 million;
  8. -borrowings from the money market will be $40 million;
  9. -non-deposit service fees should total $35 million;
  10. -a dividend payment to bank stockholders of $110 million is scheduled
Answer #1

LiquidAssets= 205+20+40 = 265

current Liab= 51+10+35+110= 216

Net Liquidity =265-216 = 49 Million

Deposit should not be considered in Current Liab

Loan Repayment should not be considered in current